Gas Y effuses half as fast as O2. What is the molar mass of gas Y? answer in g/mol
1
Expert's answer
2019-12-09T08:05:54-0500

According to Graham's law,

(r1/r2)2=M2/M1

r1 and r2 are the rates of effusion and M1 and M2 are the molar masses of the two gases. M1 = 32 g/mol

Then,

(1/2)2= 32/M of y

1/4=32/M y

M y = 32•4 = 128 g/mol


Have a nice day :)